# The Holographic QCD Running Coupling Constant from the Ricci Flow

### v1

@@ Abstract
Through a holographic model of QCD, we present a phenomenological approach to study the running of the strong coupling constant $\alpha_{s}$ in both non-perturbative and perturbative regimes. The renormalization of the metric tensor, driven by the Ricci Flow, and the breaking of conformal and chiral symmetries -thanks to introducing a double dilaton model and large-$N_{c}$ corrections- allow us to relate the existence of an infrared fixed point in the coupling constant with a smooth matching to pQCD well above 2 GeV. The proposed dilaton model yields linear Regge trajectories and vector decay constants similar to their experimental counterparts.
@@ 1 Introduction
In this work, we are interested in obtaining the running coupling constant in the strongly coupled regime and extending it to the well-known perturbative one. We can get a QCD-like running from dual models like Einstein-dilaton gravity [9, 10], or by considering Light-Front QCD embedded in a AdS background [11]. In our case we will explore the QCD running coupling constant from a different procedure: we will relate it to the running of the metric tensor on the other side of the duality. Intuitively, an abrupt change of the running coupling constant at $\Lambda_{QCD}$ could be related to an expansion of spacetime on the holographic side, which can be thought of as an exponential ansatz for the running of the metric tensor driven by the so-called Ricci Flow. The procedure above provides for a family of dilaton fields which beyond returning a running strong-coupling constant will, as a byproduct, return masses and decay constants of vector mesons.
@@ 2 Ricci Flow
where $\lambda=g_{\text{YM}}^{2}N_{c}$. We are assuming $\lambda\gg 1$, which is equivalent to supposing we are considering the large-$N_{c}$ limit in the quantum field theory at the boundary of AdS. Our interest is in QCD so $g_{\text{YM}}$ is the QCD strong coupling. In this way since $\alpha_{s}=g_{\text{YM}}^{2}/(4\pi)$, from Eq.(2) we can write:
@@
If the boundary theory has exact conformal symmetry then $\alpha_{s}$ is constant and the background remains a pure AdS spacetime. If we want to break conformal symmetry, we can consider a deviation from pure AdS imposing $R=R(z)$. In this way, $\alpha_{s}$ will depend on the holographic coordinate. More precisely, if we return to Eq.(1) and write it in the form
@@
then $g_{\mu\nu}=(R^{2}/z^{2})\eta_{\mu\nu}$ where $R^{2}\sim\sqrt{\alpha_{s}}$, see Eq.(3). We can now relate a running of $\alpha_{s}$ with a running of $g_{\mu\nu}$. In terms of $\mu=z^{-1}$, the renormalization group equation for the metric tensor $g_{\mu\nu}(\mu)=\mu^{2}R^{2}(\mu)\eta_{\mu\nu}$, with $R^{2}(\mu)=\alpha^{\prime}\sqrt{4\pi N_{c}\alpha_{s}(\mu^{2})}\eta_{\mu\nu}$, will be:
@@
In this way, we can identify the second term $\beta_{\mu\nu}=g_{\mu\nu}\frac{\mu^{2}}{4}\frac{d}{d\mu}\log\alpha_{s}$ as an homologous to a $\beta$-function of the metric $g_{\mu\nu}$, and relate naturally the running of $\alpha_{s}(\mu^{2})$ with the running of the metric tensor $g_{\mu\nu}(\mu)$. This second running is, at first order in $\alpha^{\prime}$, $\beta_{\mu\nu}=-2R_{\mu\nu}$, where $R_{\mu\nu}$ is the Ricci tensor of the metric $g_{\mu\nu}$ [12]. This result is specific for the metric Eq.(4), but motivates the introduction of the running of a general metric tensor as:
@@
where $t=\log{\mu}$. This formula is known as the Ricci Flow. Introduced in the context of string theory [12], and in mathematics [13, 14], Ricci Flow has been a method to compute the renormalization of non-linear sigma models via the above equation. Also, we can conceive the running of the metric tensor as the radial evolution equation in Hamilton-Jacobi formulation (see [15] for more details about this connection).
@@
As observed in the introduction, an abrupt change of $\alpha_{s}$ at low energies could be related to an expansion in AdS spacetime, cf. Eq.(3), since growth in $\alpha_{s}$ implies an increase in AdS radius $R$. We assume an exponential varying function of the form $\Lambda(t)=c_{0}e^{c_{1}t}$, where $c_{0}$ has units of energy squared and $c_{1}$ is unitless. We softly break the conformal symmetry by allowing $t$ dependence in $\Lambda$. The exponential parameterization will enable us to perform the desired abrupt change of $\alpha_{s}$ at low energies. Other models of conformal symmetry breaking are possible as, for example, we could model $\Lambda(t)$ as a step function to represent Hard Wall models.
@@
We fix $c_{0}=2/5\text{ GeV}^{2}$ to reproduce the usual volume form, and $c_{1}=-2$ to obtain the correct $z^{-2}$ term. We obtain:
$g_{\mu\nu}(z)=\frac{e^{\frac{2}{5}z^{2}+k}}{z^{2}}\eta_{\mu\nu}.$ (14)
The interpretation is clear: $\Lambda=\Lambda(t)$ leads to the presence of a dilaton $\phi(z)$, the field in charge of conformal symmetry breaking in the holographic model [6] as deformation of the metric tensor. An exponential ansatz is enough to find a Soft Wall-like dilaton plus a constant. The integration constant $k$ could be interpreted as an AdS radius $R_{0}^{2}$, giving a warp factor of the form $R_{0}^{2}e^{-(2/5)\phi(z)}/z^{2}$ if $k=2\log{R_{0}}$, but for now, let us consider it part of the dilaton $\phi$.
@@ 3 Double Dilaton Soft Wall model and its running
The evolution allows us to consider two different dilatons, allowing choosing a dilaton $\phi$ in $AdS$ and its opposite $-\phi$ in $AdS^{\prime}$. This setup enables us to work with a ten-dimensional model in $AdS\times AdS^{\prime}$ with action:
@@
$\phi^{\Delta}(z)=\log\left(2\cosh(\lambda^{2}(z^{2}+1))\right)\,,$ (23)
[…]
@@
To do that, and at leading order in $1/N_{c}$, let us use the expression deduced from the AdS/CFT dictionary formula seen in Eq.(16). As usual we identify $\mu\sim Q$ so we can define $\alpha_{s}(Q)=\alpha_{s}(1/z)$. Using Eq.(23) in Eq.(16) with $N_{c}=3$, and recalling experimental data is given in terms of $\alpha_{s}(Q)$, the DDSW model predicts:
@@
[…]
Even though the fit is reasonably good, we do not find pQCD running coupling constant at high energies. This is expected since holographic models of QCD work only in the strongly coupled regime. At higher energies, the running should have a transition to the well-known perturbative one. Nonetheless, this deviation occurs at higher energies compared to, for example, the Light-Front QCD AdS model employed in [11], which found a transition between the AdS and the pQCD running at around 1 GeV. The same authors improve it by joining their model with the fit parameterization to the effective strong coupling data from Ref.[25].
@@
Here $t=\alpha^{\prime}/R^{2}$, and $\chi=2-2g-L$ is the Euler characteristic that classifies Feynman diagrams in planar and non-planar diagrams by its embedding on closed surfaces [26]. $g$ is the genus of the surface and $L$ is the number of quark loops (number of boundaries). For example, planar diagrams correspond to $g=L=0$ and can be embedded in the sphere ($\chi=2$), while $g>0$ correspond to non-planar diagrams and are embedded in surfaces with holes ($\chi<2$). This classification allows $\chi$ to run from $2$ to $\infty$- ∞, including odd values because of the boundaries. The Ricci Flow is first order in $\alpha^{\prime}$ so a factor $f_{g}\left(t^{2}\right)$ is expected in the expansion. This expansion can be understood as quantum corrections to the AdS/CFT dictionary on Eq.(3) as [27]:
@@
where $R_{0}$ is the usual AdS radius. We choose the above form to be able to perform a resummation. The resummation of pQCD running coupling constant returns a Landau pole at low energies. Since Eq.(3) relates $\alpha_{s}$ with AdS geometry, a Landau pole in $\alpha_{s}$ at IR would imply a divergent metric tensor at finite $z>0$ in the UV, which is absent from our formulation. Therefore, the observed finitude of our background metric tensor is fundamental to establishing that $\alpha_{s}$ at low energies is finite. Imposing a finite metric tensor in the UV returns an IR fixed point in the running of $\alpha_{s}$. In practice, the first term of the above sum will avoid the Landau pole to find a smooth match with pQCD running.
@@
Figure 1: Model from Eq.(24) (dotted blue) and matching to perturbative QCD running coupling using DDSW model Eq.(28) for $n=2$ (dashed purple) and resumed (red), Eq.(31). The dashed vertical lines indicate the matching point with perturbative QCD for $n=2$ and the resummed version (so for $Q>2.39$ and $Q>3.79$, respectively, we have perturbative QCD running coupling in the same color).
Now, observe that the term $N_{c}^{2-b}$ can be resumed and be considered as a normalization constant $c$ for $N_{c}=3$. With all these considerations, our final parametrization is the following partial sum:
@@
Here we have included the term $N_{c}^{-2g}$ inside $a^{g}$. Thanks to the $g=0$ term, which allows a horizontal asymptote, we can match to perturbative QCD running at four loops with reference scale $M_{Z}=91.18\text{ GeV}$, $\alpha_{s}(M_{z})=0.118$. The parameters $a,b,c$ must be determined either from data or from matching to pQCD or a combination of both. In the following, we opt for matching only the parameter $c$ and let the data tell us about $a$ and $b$. To determine $c$, we can search for a matching energy point $Q_{0}$ in which pQCD running coupling, coming from high energies, meets the model of Eq.(28), coming from low energies, with the following iterative procedure. However, the matching energy point $Q_{0}$ is not known, which implies we must solve a system of two equations:
@@
In this case, and with the fitted coefficients shown in Table 2, the matching point is moved up to $Q=3.79\text{ GeV}$, see Fig.1 and Table 1. Interestingly, the resummation has a slower decrease at around $1-2$ GeV, allowing a better fit to the data.
@@
Table 1: Fit of parameters $a$ and $b$ for a fixed number of terms from 00 up to $n-1$ in the running of Eq.(28) and results for the $c$ parameter and matching point $Q_{0}$. The last row corresponds to the resummation of Eq.(31).

### v2

# Holographic QCD Running Coupling Constant from the Ricci Flow

@@ Abstract
Through a holographic model of QCD, we present a phenomenological approach to study the running of the strong coupling constant $\alpha_{s}$ in both non-perturbative and perturbative regimes. The renormalization of the metric tensor, driven by the Ricci Flow, and the breaking of conformal and chiral symmetries -thanks to introducing a double dilaton model and large-$N_{c}$ corrections- allow us to relate the existence of an infrared fixed point in the coupling constant with a smooth matching to pQCD well above 2 GeV. This is done through a model with two fit parameters and one matching point. The proposed dilaton model yields linear Regge trajectories and decay constants for scalar, vector, and tensor meson families similar to their experimental counterparts.
@@ 1 Introduction
In this work, we are interested in obtaining the running coupling constant in the strongly coupled regime and extending it to the well-known perturbative one. We can get a QCD-like running from dual models like Einstein-dilaton gravity [9, 10], or by considering Light-Front QCD embedded in a AdS background [11]. In our case we will explore the QCD running coupling constant from a different procedure: we will relate it to the running of the metric tensor on the other side of the duality. Intuitively, an abrupt change of the running coupling constant at $\Lambda_{QCD}$ could be related to an expansion of spacetime on the holographic side, which can be thought of as an exponential ansatz for the running of the metric tensor driven by the so-called Ricci Flow ([12], [13] and [14]). The procedure above provides for a family of dilaton fields which beyond returning a running strong-coupling constant will, as a byproduct, return masses and decay constants of vector mesons. When comparing to experimental data points, our objective is then to obtain a well-behaved strong coupling constant at all energy scales containing IR and UV regimes information, rather than predict them.
@@ 2 Ricci Flow
where $\lambda=g_{\text{YM}}^{2}N_{c}$. We are considering the large-$N_{c}$ limit in the quantum field theory at the boundary of AdS, so we assume $\lambda\gg 1$. Our interest is in QCD so $g_{\text{YM}}$ is the QCD strong coupling. In this way, since $\alpha_{s}=g_{\text{YM}}^{2}/(4\pi)$, from Eq.(2) we can write:
@@
If the boundary theory has exact conformal symmetry then $\alpha_{s}$ is constant and the background remains a pure AdS spacetime. If we want to break conformal symmetry, we can consider a deviation from pure AdS imposing $R=R(z)$, thus considering Eq.(2) holds approximately. In this way, $\alpha_{s}$ will depend on the holographic coordinate. More precisely, if we return to Eq.(1) and write it in the form
@@
then $g_{\mu\nu}=(R^{2}/z^{2})\eta_{\mu\nu}$ where $R^{2}\sim\sqrt{\alpha_{s}}$, see Eq.(3). We can now relate a running of $\alpha_{s}$ with a running of $g_{\mu\nu}$. In terms of $\mu=z^{-1}$, the renormalization group equation for the metric tensor $g_{\mu\nu}(\mu)=\mu^{2}R^{2}(\mu)\eta_{\mu\nu}$, with $R^{2}(\mu)=\alpha^{\prime}\sqrt{4\pi N_{c}\alpha_{s}(\mu^{2})}$, will be:
@@
In this way, we can identify the second term $\beta_{\mu\nu}=g_{\mu\nu}\frac{\mu}{4}\frac{d}{d\mu}\log\alpha_{s}$ as an homologous to a $\beta$-function of the metric $g_{\mu\nu}$, and relate naturally the running of $\alpha_{s}(\mu^{2})$ with the running of the metric tensor $g_{\mu\nu}(\mu)$. This second running is, at first order in $\alpha^{\prime}$, $\beta_{\mu\nu}=-2R_{\mu\nu}$, where $R_{\mu\nu}$ is the Ricci tensor of the metric $g_{\mu\nu}$ [12]. This result is specific for the metric Eq.(4), but motivates the introduction of the running of a general metric tensor as:
@@
where $t=\log{\mu}$. This formula is known as the Ricci Flow. Introduced in the context of string theory [12], and in mathematics [13, 14], Ricci Flow has been a method to compute the renormalization of non-linear sigma models via the above equation. Also, we can conceive the running of the metric tensor as the radial evolution equation in Hamilton-Jacobi formulation (see [15] for more details about this connection). In this direction we will treat the holographic coordinate as the evolution parameter of the flow.
@@
As observed in the introduction, an abrupt change of $\alpha_{s}$ at low energies could be related to an expansion in AdS spacetime, cf. Eq.(3), since growth in $\alpha_{s}$ implies an increase in AdS radius $R$. The function $\Lambda(t)$ will ultimately determine the warp factor of the metric tensor of the model and the running of $\alpha_{s}$. From all the possibilities we have chosen the one that induces the fastest growth yet keeping a simple ansatz, to obtain an abrupt change of $\alpha_{s}$ at low energies. Therefore, for simplicity, we assume an exponential varying function of the form $\Lambda(t)=c_{0}e^{c_{1}t}$, where $c_{0}$ has units of energy squared and $c_{1}$ is unitless. Other possibilities such as a polynomial ansatz, would induce a slower growth. We softly break the conformal symmetry by allowing $t$ dependence in $\Lambda$. The exponential parameterization will enable us to perform the desired abrupt change of $\alpha_{s}$ at low energies. Other models of conformal symmetry breaking are possible as, for example, we could model $\Lambda(t)$ as a step function to represent Hard Wall models.
@@
To fix the constants $c_{0}$ and $c_{1}$, we notice that the volume form present in the Soft Wall action is typically $\exp(-z^{2})/z^{5}$ [6]. Therefore, we fix $c_{0}=2/5\text{ GeV}^{2}$ to reproduce that usual volume form, and $c_{1}=-2$ to obtain the correct $z^{-2}$ term. We obtain:
$g_{\mu\nu}(z)=\frac{e^{c_{0}z^{2}+k}}{z^{2}}\eta_{\mu\nu}.$ (14)
This kind of background improving the original Soft Wall model have been used recently in [16], [17], [18] and [19] for different kind of purposes. In our case, we require it in order to study the strong coupling constant $\alpha_{s}$. The interpretation is clear: $\Lambda=\Lambda(t)$ leads to the presence of a dilaton $\phi(z)$, the field in charge of conformal symmetry breaking in the holographic model [6] as deformation of the metric tensor. An exponential ansatz is enough to find a Soft Wall-like dilaton plus a constant. The integration constant $k$ could be interpreted as an AdS radius $R_{0}^{2}$, giving a warp factor of the form $R_{0}^{2}e^{-(2/5)\phi(z)}/z^{2}$ if $k=2\log{R_{0}}$, but for now, let us consider it part of the dilaton $\phi$.
@@ 3 Double Dilaton Soft Wall model and its running
The evolution allows us to consider two different dilatons, allowing choosing a dilaton $\phi$ in $AdS$ and its opposite $\phi$ in $AdS^{\prime}$. This setup enables us to work with a ten-dimensional model in $AdS\times AdS^{\prime}$ with action:
@@
$\phi^{\Delta}(z)=\log\left(2\cosh(\lambda^{2}(z^{2}+k))\right)\,$, (23)
where we have used the identity $\cosh(x)=(1/2)(e^{+x}+e^{-x})$. The constant $\lambda$ is introduced to render the argument of $\cosh$ unitless, we have identified $k=k^{\prime}$ and we have chosen $k=1\text{ GeV}^{-2}$ for simplicity. To verify $\int_{0}^{\infty}e^{-\phi^{\Delta}(z)}dz<\infty$ we consider a positive dilaton. From now on we will refer to this model as the Double Dilaton Soft Wall model (DDSW).
@@
To do that, and at leading order in $1/N_{c}$, let us use the expression deduced from the AdS/CFT dictionary formula seen in Eq.(16). As usual we identify $\mu\sim Q$ so we can define $\alpha_{s}(Q)\sim\alpha_{s}(1/z)$. Using Eq.(23) in Eq.(16) with $N_{c}=3$, and recalling experimental data is given in terms of $\alpha_{s}(Q)$, the DDSW model predicts:
@@
To reach Eq.(24), we do not only embed Eq.(23) in Eq.(16) but also redefine the constants that normalize the function and render the $\cosh$ function unitless. As such, the new parameter $b$ has dimensions $\text{energy}^{-2}$], and $\lambda$ has dimensions $\text{energy}$]. Being the constant introduced in Eq.(23), $\lambda$ is analogous to the energy scale one introduces as a constant in the quadratic dilaton that defines the usual SW model. In this way, $\lambda$ should determine the mass spectrum of the model, while $b$ remains as a fit parameter in order to preserve the correct units in the cosh function.
The parameters $a$ and $b$ are determined after data fitting. We use data for the so-called effective strong coupling or effective charge $\alpha_{g1}(Q)/\pi$ collected in Ref.[27]. The effective strong coupling constant is related to $\alpha_{s}$ from the perturbative series of an observable truncated to its first order in $\alpha_{s}$ [28]. Using first-order pQCD equations makes the effective coupling renormalization-scheme and gauge independent, and free of divergence at low $Q^{2}$. Data in Ref.[27] are obtained from JLab experiments Hall CLAS EG4 (from Q = 0.143 GeV to 0.704 GeV), CLAS EG4/E977110 (from Q = 0.187 GeV to 0.490 GeV) and EG1dvcs (from $Q=0.775\text{ GeV}$ to $2.177\text{ GeV}$), represented in Fig.1 as solid star, solid circle, and solid triangle, respectively. Notice the normalization at the fixed-point $\alpha_{s}(Q\simeq 0)/\pi\simeq 1$ in Fig.1. As stated, these data extract the strong coupling constant in the effective charge approach, offering a non-perturbative renormalization-scale invariant treatment of $\alpha_{s}$ at low energies. From a fit to data with the model in Eq.(24), we obtain $a=1.545\pm 0.047$ $b=1.150\pm 0.041\text{ GeV}^{-2}$ with a reasonable $\chi^{2}/DOF=1.45$. As can be seen in Fig.1, we obtain naturally an infrared fixed point. At $Q\simeq 1.5\text{ GeV}$ we observe a slight deviation from the experimental result and a tendency for $\alpha_{s}(Q)/\pi$ to go to zero quite fast. We have also compared in Fig.1 with the holographic light-front model ([11], $\alpha^{AdS}_{Modified,g1}$ Eq.(12)). The base holographic light-front result is not displayed, being very similar to the DDSW result in blue in Fig.1.
Even though the fit is reasonably good, we do not find pQCD running coupling constant at high energies. This is expected since holographic models of QCD work only in the strongly coupled regime. At higher energies, the running should have a transition to the well-known perturbative one. Nonetheless, this deviation occurs at higher energies compared to, for example, the Light-Front QCD AdS model employed in [11], which found a transition between the AdS and the pQCD running at around 1 GeV. The same authors improve it by joining their model with the pQCD prediction, resulting in a better result in the perturbative regime.
@@
Here $t=\alpha^{\prime}/R^{2}$, and $\chi=2-2g-L$ is the Euler characteristic that classifies Feynman diagrams in planar and non-planar diagrams by its embedding on closed surfaces [30]. $g$ is the genus of the surface and $L$ is the number of quark loops (number of boundaries). For example, planar diagrams correspond to $g=L=0$ and can be embedded in the sphere ($\chi=2$), while $g>0$ correspond to non-planar diagrams and are embedded in surfaces with holes ($\chi<2$). This classification allows $\chi$ to run from $2$ to -$\infty$, including odd values because of the boundaries. The Ricci Flow is first order in $\alpha^{\prime}$ so a factor $f_{g}\left(t^{2}\right)$ is expected in the expansion. This expansion can be understood as quantum corrections to the AdS/CFT dictionary on Eq.(3) as [31]:
@@
where $R_{0}$ is the usual AdS radius. We choose the above form to be able to perform a resummation. The expansion parameter we use here is $1/N_{c}$, considering $N_{c}\to\infty$ grows faster than $R^{2}/\alpha^{\prime}$. The resummation of pQCD running coupling constant returns a Landau pole at low energies. Since Eq.(3) relates $\alpha_{s}$ with AdS geometry, a Landau pole in $\alpha_{s}$ at IR would imply a divergent metric tensor at finite $z>0$ in the UV, which is absent from our formulation. Therefore, the observed finitude of our background metric tensor is fundamental to establishing that $\alpha_{s}$ at low energies is finite. Imposing a finite metric tensor in the UV returns an IR fixed point in the running of $\alpha_{s}$. In practice, the first term of the above sum will avoid the Landau pole to find a smooth match with pQCD running.
@@
Figure 1: Model from Eq.(24) (dotted blue), matching to perturbative QCD running coupling using DDSW model Eq.(28) for $n=2$ (dashed purple), resummed (red) Eq.(31). Their $\chi^{2}/DOF$ values are $1.45$, $0.99$ and $0.93$, respectively. We also compare with the holographic light-front model [11] (dotted black). The dashed vertical lines indicate the matching point with perturbative QCD for $n=2$ and the resummed version (so for $Q>2.39$ and $Q>3.79$, respectively, we have perturbative QCD running coupling in the same color).
Now, observe that the term $N_{c}^{2-b}$ can be resummed and be considered as a normalization constant $c$ for $N_{c}=3$. With all these considerations, our final parametrization is the following partial sum:
@@
Here we have included the term $N_{c}^{-2g}$ inside $a^{g}$. Thanks to the $g=0$ term, which allows a horizontal asymptote, we can match to perturbative QCD running at four loops with reference scale $M_{Z}=91.18\text{ GeV}$, $\alpha_{s}(M_{z})=0.118$. The parameters $a,b,c$ must be determined either from data or from matching to pQCD or a combination of both. Having three parameters to adjust, we acknowledge the limitations of the model (28). We believe the model could be improved in future works to have more predictive power. In the following, we opt for matching only the parameter $c$ and let the data tell us about $a$ and $b$. By matching, we mean we normalize at $Q_{0}$ both image and derivative of the non-perturbative and perturbative results for the strong coupling. In this way we determine the $c$ parameter solving a system of equations, while a fit to the data will determine $a$ and $b$ parameters. To determine $c$, we can search for a matching energy point $Q_{0}$ in which pQCD running coupling, coming from high energies, meets the model of Eq.(28), coming from low energies, with the following iterative procedure. However, the matching energy point $Q_{0}$ is not known, which implies we must solve a system of two equations:
@@
In this case, and with the fitted coefficients shown in Table 2, the matching point is moved up to $Q=3.79\text{ GeV}$, see Fig.1 and Table 1. Interestingly, the resummation has a slower decrease at around 121-2 GeV, allowing a better fit to the data.
@@
Table 1: Fit of parameters $a$ and $b$ for a fixed number of terms from 0 up to n1n-1 in the running of Eq.(28) and results for the $c$ parameter and matching point $Q_{0}$. The last row corresponds to the resummation of Eq.(31).